P152

IR STUDY OF METALLOTROPIC REARRANGEMENTS AND ION-PAIR INTERACTIONS IN CHROMIUM TRICARBONYL COMPLEXES WITH POLYCYCLIC AROMATIC LIGANDS

\underline{B. V. LOKSHIN}, M. G. EZERNITSKAYA, V. A. MALTZEVA, A.N.Nesmeyanov Institute of Organoelement Compounds, Russian Academy of Sciences, 28 Vavilova Str., 117813 Moscow, Russia; Yu. F. OPRUNENKO and Yu. A. USTYNYUK, Department of Chemistry, M.V.Lomonosov Moscow State University, 119899, Vorob’evy Gory, Moscow, Russia.

Metallotropic rearrangements (itramolecular rearrangements involving the displacement of the transition metal atom) in chromium tricarbonyl π-complexes with polycyclic aromatic ligands were studied by FTIR spectroscopy. A novel η^{6} η^{5} rearrangements have been found and studied for anions η^{6}-phenalenyl- (I) and η^{6}-(9- hydroanthracenyl) (II) tricarbonychromium generated by deprotonation of corresponding neutral η^{6}-complexes with various bases (BuLi, Ph_{3}P=CH_{2}, proton sponges or t-BuOM, where M= K, Cs) at low temperatures. IR spectra in the ν(CO) range of the η^{6}- and η^{5}-anions with various counter-ions (Li^{+}, K^{+}, and Cs^{+}) in THF and dimethoxyethane solutions and in the presence of the solvating additive, dibenzo-18-crown-6, in a temperature region between -70 to 20^{o}C showed that anionic complexes exist in solution as tight and solvent separated ion pairs or as their equilibrium mixture. The factors that influence the ion-pair equilibrium (solvent, the excess of the base, temperature, the nature of the counter-ion) have been studied. The results are compared with the NMR data.. The work was supported by the Russian Foundation for Basic Research (grants 96-03-3241a and 96-03-34082a).